Assume that The AM Bakery is preparing a budget for the month ending October 31. Management prepares the budget by starting with the actual results for August 31. Next, management considers what the differences in costs will be between August and October.

Management expects revenue in October to be 10 percent more than in August, and it expects all ingredient costs (e.g., flour, butter, and so on) to be 10 percent higher in October than in August. Management expects “other” labor costs to be 20 percent higher in October than in August, partly because more labor will be required in October and partly because employees will receive a pay raise. The manager will receive a pay raise that will increase his salary from $6,200 in August to $6,870 in October. Rent, utilities, and marketing costs are not expected to change.

Required:

Prepare a budget for The AM Bakery for October.

THE AM BAKERY
Bakery Sales
Budgeted Costs
For the Month Ending October 31
Actual Budgeted
(August) (October)
Ingredients
Flour $5,600
Butter 5,200
Oil 3,400   
Fruit 3,000
Nuts 2,600
Chocolate 1,650
Other 1,250
Total ingredients $22,700
Labor
Channel manager $6,200
Other 12,400
Utilities 4,100
Rent 5,300
Marketing 1050
Total bakery cost $51,750
Revenues 69,200
Answer #1
THE AM BAKERY
Bakery Sales
Budgeted Costs
For the Month Ending October 31
Actual Budgeted
(August) (October)
Ingredients
Flour 5600 6160 =5600*1.1
Butter 5200 5720 =5200*1.1
Oil 3400 3740 =3400*1.1
Fruit 3000 3300 =3000*1.1
Nuts 2600 2860 =2600*1.1
Chocolate 1650 1815 =1650*1.1
Other 1250 1375 =1250*1.1
Total ingredients 22700 24970
Labor
Channel manager 6200 6870
Other 12400 14880 =12400*1.2
Utilities 4100 4100
Rent 5300 5300
Marketing 1050 1050
Total bakery cost 51750 57170
Revenues 69200 76120 =69200*1.1